the sum of the ages of a mother & her daughter is 60 years.12 years ago the mother was eight times as old as her daughter.how old is the daughter now?
Answer by malglu(63)

let age mum now be m
age daughter now =d
1)m+d=60
2)m-12 = 8(d-12)
m-12 =8d -96
rearrange to make
m=8d-84
substitute this into the first equation to make
8d-84+d=60
rearrange and solve
9d-84=60
9d=144
d=16